A baseball mass 250.00 g approaches a bat horizontally at a speed of 40.20 m/s and is hit straight back at a speed of 51.50 m/s . If the ball is in contact with the bat for a time of 1.60 ms, what is the average force exerted on the ball by the bat? Neglect the weight of the ball, since it is so much less than the force of the bat. Choose the direction of the incoming ball as the positive direction.
Answer #1

By using the impulse momentum,

P = F*t

F = P/t

= (mVf - mVi) /t

= m*(Vf - Vi) /t

= 0.25*(-51.50 - 40.20)/(1.6*10^-3)

= - 14328.125 N

= - 14.328 kN

By using Newtons 3rd law, we can say that the force of the ball on the bat is equal to the force of the bat on the ball.

So, the direction of the force is only a matter of perspective.

Because of Newton’s 3rd law, the force of the ball on the bat is equal to the force of the bat on the ball. The direction of the force is only a matter of perspective.